# Measurement of medium–induced acoplanarity in central $\mathrm{Au}\text{--}\mathrm{Au}$ and $pp$ collisions at $\sqrt{s_{\mathrm{NN}}}=200$ GeV using direct–photon+jet and $\pi^{0}$+jet correlations

### v1

@@ Abstract
Figure 3: Left panels: ${I}_{\mathrm{AA}}(\Delta\phi)$ distributions for $1511<E_{\mathrm{T}}^{\mathrm{trig}}<$ 15 GeV with recoil jet $R=0.2$ and 0.5. Top: $\gamma_{\mathrm{dir}}$ trigger, $1510<p_{\mathrm{T,jet}}^{\mathrm{ch}}<$ 15 $\mathrm{GeV/}c$; middle: $\pi^{0}$  trigger, $1510<p_{\mathrm{T,jet}}^{\mathrm{ch}}<$ 15 $\mathrm{GeV/}c$; bottom: $\pi^{0}$ trigger, $2015<p_{\mathrm{T,jet}}^{\mathrm{ch}}<$ 20 $\mathrm{GeV/}c$. Arrows indicate 95% CL. Theoretical calculations are discussed in the text. Right panels: integral of ${I}_{\mathrm{AA}}(\Delta\phi)$ over $2.35<\Delta\phi<\pi$ [33].
@@
For large acoplanarity, all panels show suppression of ${I}_{\mathrm{AA}}(\Delta\phi)$ for recoil jets with $R=0.2$ and significant enhancement for $R=0.5$. The value of ${I}_{\mathrm{AA}}(\Delta\phi)$ at $\Delta\phi\approx 2.65$ differs for $R=0.2$ and 0.5 by a factor $20\pm 2$ (sys) for $\pi^{0}$ triggers and recoil jet $2015<p_{\mathrm{T,jet}}^{\mathrm{ch}}<$ 20 $\mathrm{GeV/}c$ (bottom panel), and a factor $3.8\pm 1.7$ for $\gamma_{\mathrm{dir}}$ triggers and recoil jet $1510<p_{\mathrm{T,jet}}^{\mathrm{ch}}<$ 15 $\mathrm{GeV/}c$ (top panel); statistical error is negligible. Significant differences are likewise observed for $\Delta\phi\approx 2.35$. A similar, marked $R$–dependent broadening of ${I}_{\mathrm{AA}}(\Delta\phi)$ was observed in the same $p_{\mathrm{T,jet}}^{\mathrm{ch}}$-range for h+jet correlations in central $\mathrm{Pb}\text{--}\mathrm{Pb}$ collisions at $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV [27].
Medium–induced yield enhancement at large acoplanarity may arise from secondary partonic scattering with QGP quasiparticles [19, 20, 21, 22, 23, 24]. However, such scattering effects should be evident for all $R$–values, which are used to probe the population of hard-scattering processes with different apertures. In contrast, Fig. 3 and Ref. [27] show selective enhancement for $R=0.5$ but not $R=0.2$, which is not consistent with jet scattering as the predominant underlying mechanism.
@@
JEWEL is based on PYTHIA, which describes the $pp$ measurements well (Fig. 2). JEWEL also describes well the $R$–dependent medium–induced acoplanarity broadening seen at the LHC [27]. However, at RHIC energies JEWEL does not exhibit significant medium–induced broadening for either $R=0.2$ and 0.5, for both $\gamma_{\mathrm{dir}}$ and $\pi^{0}$ triggers (Fig. 3).
[…]
@@
The Hybrid Model predicts medium–induced narrowing for both $R=0.2$ and $R=0.5$ (Fig. 3), in disagreement with the data, with similar predictions at LHC energies [27]. The End Matter presents Hybrid Model calculations for recoil jet $105<p_{\mathrm{T,jet}}^{\mathrm{ch}}<$ 10 $\mathrm{GeV/}c$ which exhibit $R$–dependent broadening whose qualitative features are similar to those seen in data for larger values of $p_{\mathrm{T,jet}}^{\mathrm{ch}}$, but only with wake implemented. This provides additional insight into the physical origin of the broadening, and its modeling.
@@
Summary- This Letter reports measurements of the semi–inclusive acoplanarity distribution of charged-particle jets recoiling from $\gamma_{\mathrm{dir}}$ and $\pi^{0}$ triggers in $pp$ and central $\mathrm{Au}\text{--}\mathrm{Au}$ collisions at $\sqrt{s_{\mathrm{NN}}}=200$ GeV. Significant $R$–dependent medium–induced acoplanarity broadening is observed, corresponding to a yield enhancement for large $R$ compared to small $R$ jets of up to a factor 20. A picture that accommodated these observations and a corresponding measurement at the LHC is that the broadening arises predominantly from diffuse QGP medium response to the passage of an energetic parton, i.e. the jet wake, rather than single hard Rutherford–like scattering off of QGP quasiparticles. Theoretical calculations incorporating jet quenching and QGP wake effects exhibit significant differences with the measurements, requiring modification of their underlying physics description to improve the agreement. These measurements provide new insight into the nature of the interaction between jets and the QGP, and the application of jets to probe QGP dynamics.
@@
Acknowledgments- We thank Jaime Norman, Yu Shi, Shu-Yi Wei, Bowen Xiao, Feng Yuan, Danny Pablos, and Krishna Rajagopal for providing calculations. We thank the RHIC Operations Group and SDCC at BNL, the NERSC Center at LBNL, and the Open Science Grid consortium for providing resources and support. This work was supported in part by the Office of Nuclear Physics within the U.S. DOE Office of Science, the U.S. National Science Foundation, National Natural Science Foundation of China, Chinese Academy of Science, the Ministry of Science and Technology of China and the Chinese Ministry of Education, NSTC Taipei, the National Research Foundation of Korea, Czech Science Foundation and Ministry of Education, Youth and Sports of the Czech Republic, Hungarian National Research, Development and Innovation Office, New National Excellency Programme of the Hungarian Ministry of Human Capacities, Department of Atomic Energy and Department of Science and Technology of the Government of India, the National Science Centre and WUTID-UB of Poland, the Ministry of Science, Education and Sports of the Republic of Croatia, German Bundesministerium für Bildung, Wissenschaft, Forschung and Technologie (BMBF), Helmholtz Association, Ministry of Education, Culture, Sports, Science, and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) and Agencia Nacional de Investigación y Desarrollo (ANID) of Chile.

### v2

@@ III Results
Figure 3: Left panels: ${I}_{\mathrm{AA}}(\Delta\phi)$ distributions for $11<E_{\mathrm{T}}^{\mathrm{trig}}<15$ GeV with recoil jet $R=0.2$ and 0.5. Top: $\gamma_{\mathrm{dir}}$ trigger, $10<p_{\mathrm{T,jet}}^{\mathrm{ch}}<15$ $\mathrm{GeV/}c$; middle: $\pi^{0}$ trigger, $10<p_{\mathrm{T,jet}}^{\mathrm{ch}}<15$ $\mathrm{GeV/}c$; bottom: $\pi^{0}$ trigger, $15<p_{\mathrm{T,jet}}^{\mathrm{ch}}<20$ $\mathrm{GeV/}c$. Arrows indicate 95% CL. Theoretical calculations are discussed in the text. Right panels: integral of ${I}_{\mathrm{AA}}(\Delta\phi)$ over $3\pi/4<\Delta\phi<\pi$ [5].
@@
For large acoplanarity, all panels show suppression of ${I}_{\mathrm{AA}}(\Delta\phi)$ for recoil jets with $R=0.2$ and significant enhancement for $R=0.5$. The value of ${I}_{\mathrm{AA}}(\Delta\phi)$ at $\Delta\phi\approx 2.65$ differs for $R=0.2$ and 0.5 by a factor $20\pm 2$ (sys) for $\pi^{0}$ triggers and recoil jet $15<p_{\mathrm{T,jet}}^{\mathrm{ch}}<20$ $\mathrm{GeV/}c$ (bottom panel), and a factor $3.8\pm 1.7$ for $\gamma_{\mathrm{dir}}$ triggers and recoil jet $10<p_{\mathrm{T,jet}}^{\mathrm{ch}}<15$ $\mathrm{GeV/}c$ (top panel); statistical error is negligible. Significant differences are likewise observed for $\Delta\phi\approx 2.35$. A similar, marked RR–dependent broadening of ${I}_{\mathrm{AA}}(\Delta\phi)$ was observed in the same $p_{\mathrm{T,jet}}^{\mathrm{ch}}$-range for h+jet correlations in central $\mathrm{Pb}\text{--}\mathrm{Pb}$ collisions at $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV [7].
Medium–induced yield enhancement at large acoplanarity may arise from secondary partonic scattering with QGP quasiparticles [15, 20, 35, 34, 31, 36]. However, such scattering effects should be evident for all RR–values, which are used to probe the population of hard-scattering processes with different apertures. In contrast, Fig. 3 and Ref. [7] show selective enhancement for $R=0.5$ but not $R=0.2$, which is not consistent with jet scattering as the predominant underlying mechanism.
@@
JEWEL is based on PYTHIA, which describes the $pp$ measurements well (Fig. 2). JEWEL also describes well the RR–dependent medium–induced acoplanarity broadening seen at the LHC [7]. However, at RHIC energies JEWEL does not exhibit significant medium–induced broadening for either $R=0.2$ and 0.5, for both $\gamma_{\mathrm{dir}}$ and $\pi^{0}$ triggers (Fig. 3).
For the analytic QCD calculation the Gaussian broadening width is $\langle\hat{q}{L}\rangle$, where $\hat{q}$ is the jet transport coefficient [44], $L$ is the in–medium path length, and $\langle{\ldots}\rangle$ indicates averaging over collisions. The band in Figure 3 corresponds to $3<\langle\hat{q}{L}\rangle<13$ GeV2, reproducing the measured ${I}_{\mathrm{AA}}(\Delta\phi)$ for $R=0.5$, but not for $R=0.2$. This feature is intrinsic to the simple Gaussian azimuthal broadening employed, which models jet–QGP multiple scattering without $R$ dependence. This disagreement with data provides additional evidence that the observed marked RR–dependence does not arise predominantly from in–medium soft scattering.
@@
The Hybrid Model predicts medium–induced narrowing for both $R=0.2$ and $R=0.5$ (Fig. 3), in disagreement with the data, with similar predictions at LHC energies [7]. The End Matter presents Hybrid Model calculations for recoil jet $5<p_{\mathrm{T,jet}}^{\mathrm{ch}}<10$ $\mathrm{GeV/}c$ which exhibit RR–dependent broadening whose qualitative features are similar to those seen in data for larger values of $p_{\mathrm{T,jet}}^{\mathrm{ch}}$, but only with wake implemented. This provides additional insight into the physical origin of the broadening, and its modeling.
IV Summary
@@
This article reports measurements of the semi–inclusive acoplanarity distribution of charged-particle jets recoiling from $\gamma_{\mathrm{dir}}$ and $\pi^{0}$ triggers in $pp$ and central $\mathrm{Au}\text{--}\mathrm{Au}$ collisions at $\sqrt{s_{\mathrm{NN}}}=200$ GeV. Significant RR–dependent medium–induced acoplanarity broadening is observed, corresponding to a yield enhancement for large $R$ compared to small $R$ jets of up to a factor 20. A picture that accommodated these observations and a corresponding measurement at the LHC is that the broadening arises predominantly from diffuse QGP medium response to the passage of an energetic parton, i.e. the jet wake, rather than single hard Rutherford–like scattering off of QGP quasiparticles. Theoretical calculations incorporating jet quenching and QGP wake effects exhibit significant differences with the measurements, requiring modification of their underlying physics description to improve the agreement. These measurements provide new insight into the nature of the interaction between jets and the QGP, and the application of jets to probe QGP dynamics.
Acknowledgements.
@@
We thank Jaime Norman, Yu Shi, Shu-Yi Wei, Bowen Xiao, Feng Yuan, Danny Pablos, and Krishna Rajagopal for providing calculations. We thank the RHIC Operations Group and SDCC at BNL, the NERSC Center at LBNL, and the Open Science Grid consortium for providing resources and support. This work was supported in part by the Office of Nuclear Physics within the U.S. DOE Office of Science, the U.S. National Science Foundation, National Natural Science Foundation of China, Chinese Academy of Science, the Ministry of Science and Technology of China and the Chinese Ministry of Education, NSTC Taipei, the National Research Foundation of Korea, Czech Science Foundation and Ministry of Education, Youth and Sports of the Czech Republic, Hungarian National Research, Development and Innovation Office, New National Excellency Programme of the Hungarian Ministry of Human Capacities, Department of Atomic Energy and Department of Science and Technology of the Government of India, the National Science Centre and WUT ID-UB of Poland, German Bundesministerium für Bildung, Wissenschaft, Forschung and Technologie (BMBF), Helmholtz Association, Ministry of Education, Culture, Sports, Science, and Technology (MEXT), and Japan Society for the Promotion of Science (JSPS).
Appendix A Hybrid Model comparison
Figure A1: Data from Fig. 3 for recoil jet $10<p_{\mathrm{T,jet}}^{\mathrm{ch}}<15$ $\mathrm{GeV/}c$, and for Hybrid Model calculations for both $10<p_{\mathrm{T,jet}}^{\mathrm{ch}}<15$ $\mathrm{GeV/}c$ and $5<p_{\mathrm{T,jet}}^{\mathrm{ch}}<10$ $\mathrm{GeV/}c$, both with and without wake.
Figure A1 shows the $\gamma_{\mathrm{dir}}$ and $\pi^{0}$–triggered ${I}_{\mathrm{AA}}(\Delta\phi)$ distributions for recoil jet $10<p_{\mathrm{T,jet}}^{\mathrm{ch}}<15$ $\mathrm{GeV/}c$ (Fig. 3), and for Hybrid Model calculations for $10<p_{\mathrm{T,jet}}^{\mathrm{ch}}<15$ $\mathrm{GeV/}c$ and $5<p_{\mathrm{T,jet}}^{\mathrm{ch}}<10$ $\mathrm{GeV/}c$, both with and without wake. For both triggers, Hybrid Model calculations exhibit no significant $R$-dependence without wake in both $p_{\mathrm{T,jet}}^{\mathrm{ch}}$ intervals, or with wake for $10<p_{\mathrm{T,jet}}^{\mathrm{ch}}<15$ $\mathrm{GeV/}c$. However, for $5<p_{\mathrm{T,jet}}^{\mathrm{ch}}<10$ $\mathrm{GeV/}c$, the calculations with wake exhibit a marked $R$-dependent acoplanarity broadening for both triggers. Although this is qualitatively similar to the effect seen in data for $10<p_{\mathrm{T,jet}}^{\mathrm{ch}}<15$ $\mathrm{GeV/}c$, the theoretical model shows a larger effect for $\pi^{0}$ triggers than $\gamma_{\mathrm{dir}}$ triggers.
The striking $R$-dependent medium–induced acoplanarity broadening seen in data is therefore reproduced qualitatively by the Hybrid Model, though only with the hydrodynamic wake implemented and only in a lower kinematic interval than the reported measurement.
Appendix B Variation of ME normalization region
Figure B2: Distributions of $\widetilde{Y}_{\Delta\phi}$($p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}$) from PRL Fig. [3] using the nominal ME normalization region (blue), and those where the upper limit of the normalization region are reduced by 1 (red) or 3 (green) $\mathrm{GeV/}c$.
This section documents the systematic uncertainty due to variation in the limits of the ME normalization region. Figure B2 shows the distribution of $\widetilde{Y}_{\Delta\phi}$($p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}$) (Eq. [3]) for the nominal ME normalization region (Fig. [1]) and for shifting the upper limit of the normalization region lower by 1 and 3 $\mathrm{GeV/}c$. The resulting variation in $\widetilde{Y}_{\Delta\phi}$($p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}$) is seen to be small, particularly for $p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}>0$, but is nevertheless incorporated into the systematic uncertainty of the measurement.
Appendix C Effect of $\Delta\phi$-smearing correction
Figure C3: Distributions of $\widetilde{Y}_{p_{\mathrm{T}}}$($\Delta\phi$) with bin-wise correction for $\Delta\phi$-smearing from Fig. [2] (filled boxes), and the central value of the distributions without the bin-wise correction (open circles).
This section documents the bin-wise correction for $\Delta\phi$-smearing applied to the measurement of $\widetilde{Y}_{p_{\mathrm{T}}}$($\Delta\phi$) . Figure C3 shows $\widetilde{Y}_{p_{\mathrm{T}}}$($\Delta\phi$) distributions with the bin-wise correction (Fig. [2])), and the central value of the same distributions without the bin-wise correction.
Figure C4: Distributions of ${I}_{\mathrm{AA}}(\Delta\phi)$ (Fig. [3]) (colored markers) compared to the central values of the same ratios without the bin-wise correction for $\Delta\phi$-smearing.
Figure C4 shows ${I}_{\mathrm{AA}}(\Delta\phi)$ with the bin-wise correction (Fig. [3]), and its central value without the bin-wise correction. Figures C3 and C4 show that the magnitude of the bin-wise correction for smearing in $\Delta\phi$ is sub-leading relative to other effects contributing to the systematic uncertainty, and is negligible in comparison to the magnitude of medium-induced modification observed in the measurements of ${I}_{\mathrm{AA}}(\Delta\phi)$.